OpenStudy (anonymous):

a = 40, c = 41, b = help solve pleaz

OpenStudy (aaronandyson):

\[A^2 + B^2 = C^2\]

OpenStudy (aaronandyson):

Where 'a' is 40 and 'c' is 41

OpenStudy (aaronandyson):

so to find b

OpenStudy (aaronandyson):

so to find b

OpenStudy (aaronandyson):

\[B^2 = C^2 - A^2\]

OpenStudy (aaronandyson):

So whats 41*41?

OpenStudy (aaronandyson):

@lokitroll97 ?

OpenStudy (anonymous):

yes

OpenStudy (aaronandyson):

So now solve and tell me what did you get on solving .

OpenStudy (anonymous):

1600+1681=C^2 --- put together and sqrted = 57.28

OpenStudy (anonymous):

hello

OpenStudy (sleepyjess):

\(40^2+b^2=41^2\) What is \(40^2\)?

OpenStudy (anonymous):

1600

OpenStudy (sleepyjess):

Ok, and \(41^2\)?

OpenStudy (anonymous):

41 is 1681 add and sqrt it and it = 57.28

OpenStudy (sleepyjess):

No, we aren't adding this time

OpenStudy (anonymous):

oh k -

OpenStudy (sleepyjess):

We get \(1600 + b^2 = 1681\). We need to subtract 1600 from both sides

OpenStudy (anonymous):

so 81 = b^2

OpenStudy (sleepyjess):

yes, and the sqrt81?

OpenStudy (anonymous):

ninner

OpenStudy (sleepyjess):

I am going to assume that means 9, and if so, then correct

OpenStudy (anonymous):

yes military or air force to be exact

OpenStudy (sleepyjess):

huh?

OpenStudy (anonymous):

its how a pilot communicates the number nine when communicating via telcom